a 5 percent salt solution is mixed with a 10% salt solution. How many mililiters of each solution are needed to obtain 400 mililiters of an 8% solution?
Answer by stanbon(75887)

A 5 percent salt solution is mixed with a 10% salt solution. How many mililiters of each solution are needed to obtain 400 mililiters of an 8% solution?
--------------------
Equation:
salt + salt = salt
0.05x + 0.10(400-x) = 0.08*400
Multiply thru by 100
5x + 10*400 - 10x = 8*400
-5x = -2*400
x = 2*80
x = 160 milimeters (amt. of 5% solution needed for the mixture)
---
400-160 = 340 milimeter (amt of 10% solution needed for the mixture)
=====================================================================
Cheers,
Stan H.